OpenStudy (anonymous):

The sum of the digits of a two digit number is 11 and the difference of the digit is 5. What is the number?

OpenStudy (aakashsudhakar):

x+y=11 x-y=5 Basic algebraic substitution follows. I'd recommend substituting 5-y for x (from the second equation).

OpenStudy (loser66):

apply this trick big + small = sum big - small = dif big = (sum + dif )/2 small = (sum -dif) /2